# Discovery of Volatile Gas in the Giant Impact Disk around the 150-Myr old HD 23514

### v1

@@ C.4 The Nature of the 15 µm bump in the Spitzer/IRS Spectrum
We investigated whether the 15 µm bump in the Spitzer spectrum as shown in Figure 3 could be due to data reduction artifacts by examining the CASSIS IRS spectra of a handful debris systems (e.g., HD 172555, HD 15407 and HD 69830 and others), which are brighter or at a similar brightness as HD 23514 and observed in the same observing mode, using the CASSIS products. We found no similar bump near 15 µm, suggesting that the 15 µm bump is likely astrophysical in nature.
@@
Appendix D Estimating the Dust Location, Blowout Sizes and Various Timescales
D.1 SED Models
@@
The purpose of the SED modeling is to estimate the disk extent and to illustrate that the bulk of the silica dust is co-located with the hot molecular gas (from the temperature argument). The derived parameters are then used to estimate various timescales that are relevant in HD 23514.
@@
Given the number of assumed parameters, the SED models are extremely degenerate and sensitive to the disk density distribution, and only serve as a zero-order estimate without other constraints like resolved disk images. We stress that the SED modeling results are only applicable to the part of the disk where its bulk emission is fully captured by the SED measurements. The assumptions for our SED models are: (1) the star is the only heating source and in an optically thin environment, (2) there is one single, flat (constant surface density), continuous disk with a radial range of inner () and outer () radii from the star, and (3) the dust grains are all in one uniform size distribution with a size power-law index of 3.5 and a maximum size of 1000 µm. We explored several minimum grain sizes and found that a minimum size of 0.1–0.5 µm silica- and silicate-like grains are needed in order to produce the prominent 10/20 µm features, even though the typical radiation blowout size is 1 µm given the stellar mass and luminosity (1.3 and 3 ). The dynamical stability of these small grains is subject to their exact optical properties in response to the stellar radiation, which is not explored/considered here. We consider that the impact of other parameters such as the maximum grain size () and surface density power-law is minimal in a way that (1) affects the total dust mass in a fashion under the assumed size distribution, and (2) a steeper surface density distribution would extend so enough colder material can account for the total emission.
@@
A large uncertainty in our SED model resides in the grain composition. The dust temperatures depend sensitively on the grain absorption efficiency at the wavelengths where the star emits most its energy (i.e., UV and optical for HD 23514). Sub-µm grains are generally hotter than their larger counterparts and the exact temperatures can be vastly different between different types of composition as shown in the left panel of Figure 9. Because the carbonaceous and silica grains are less absorptive than the typical silicates in the UV and optical, they generally have lower temperatures than that of silicates at the same stellocentric distance from the star.
@@
Determining detailed dust mineralogy to infer the exact dust composition is beyond the scope of this short paper. We only explored limited dust compositions that are commonly found in thermally processed dust in circumstellar environments. For the dust emission calculation, we adopted the grain properties computed by optool (Dominik et al., 2021) using the Distribution of Hollow Spheres (DHS, Min et al. 2005) approach with an irregularity factor fmax of 0.8. We found that the system’s SED can be well represented by three types of dust: silica (density = 2.65 g cm-3, Henning & Mutschke 1997), pyroxene ( = 3.2 g cm-3, Dorschner et al. 1995) and amorphous carbon ( = 1.85 g cm-3, Zubko et al. 1996) where the first two give rise to the prominent dust features in the mid-IR while the amorphous carbon and a small fraction of Mg-rich pyroxene grains contribute the underlining featureless dust continuum as shown in Figure 4. A single component disk with = 0.1 au and 2–3 au gives a good fit to the measured SED except that the silica grains need to be restricted within 0.1–0.2 au to account for enough flux in the 3–5 µm range without changing the contrast/ratio between the 10 and 20 µm features (the right panel of Figure 9). In other words, having silica grains beyond 1 au would shift the near-IR flux toward the far-IR because of its low thermal temperature outside 1 au, resulting in poor fits between 3 and 5 µm. We note that the model produces a much narrower 10 µm feature than the observed one, likely due to the mismatch in the various silica polymorphs (Koike et al., 2013).
@@
Based on the SED model, the system’s infrared fractional luminosity () is 1.76 with 70% coming from the featureless dust continuum. The total dust mass is g (1.67) and the feature-producing grains account for 5% of the mass. Because the SED models prefer a confined (sub-au) region for the silica grains, the total silica dust mass (integrated to 1 mm in size) is on the order of 1% (1022 g) of the total dust mass. In summary, the bulk of the dust in HD 23514 is within 3 au and the silica dust is mostly located inside a sub-au region, similar to the location of the hot molecular gas found in Appendix C.1. HD 23514 was not detected by the ALMA 1.3 mm shallow survey for the Pleiades (Sullivan et al., 2022), suggesting lack of cold (40 K) dust, although the ALMA data could not exclude a less dusty planetesimal belt like our current Kuiper belt. In conclusion, the HD 23514 system does not possess a massive (tens of au) Kuiper-belt-like analog that are found in 20% of solar-like stars (Sierchio et al., 2014; Matrà et al., 2025).
D.2 Radiation Pressure Blowout Sizes around HD 23514
@@
In gas-free debris disks, stellar radiation pressure is one of the most important forces acting on grains that counters with the star’s gravity and dictates the dust dynamics, which can be characterized by the parameter defined as the ratio of radiation pressure to gravitational forces, i.e.,
@@
where and are luminosity and mass of the star, is grain size, is grain density, and is the radiation-pressure efficiency averaged over stellar spectrum (Burns et al., 1979). The last three parameters have strong dependency on the exact grain properties (composition, size, density and porosity, as detailed in Arnold et al. 2019). Grains with get pushed into hyperbolic orbits and leave the system, while grains with are on elliptical orbits and spiral inward due to Poynting-Robertson (P-R) drag. The radiation pressure blowout size is defined as , and the exact size depends on the type of star, and most importantly, the grain properties as illustrated in Figure 10. The values are calculated using the stellar Kurucz model described in Appendix A.1, the grain compositions adopted in Appendix D.1 with two additions (olivine: 3.71 from Dorschner et al. 1995, and silica: 2.65 from Kitamura et al. 2007), and assuming compact (i.e., ideal with zero porosity) grain shape. For nominal silicate and carbonaceous compositions, grains small than the blowout size (1 µm for HD 23514) are ejected and on a terminal velocity of (Su et al., 2005). While for silica composition, some (such as the one with the optical constant from Henning & Mutschke 1997) could have 0.5 for all sizes, and some (from Kitamura et al. 2007) would have the 0.5 grains within two blowout sizes. This also illustrates the blowout size is not a strict number, highly depending on the exact composition. We stress the importance of grain optical properties in the UV and optical wavelengths where the star emits most of its energy in estimating the grain’s values.
@@ D.3 Relevant Timescales
For a dust-dominated debris disk, the dominant grain loss mechanisms are collisional grinding and P-R/stellar wind drags. Once the grains generated by collisional cascades are smaller than the radiation blowout size (1 µm in HD 23514 when the ratio between the radiation pressure force and gravity, defined as , equal to 0.5), they are ejected from the system by radiation pressure under normal conditions. The blowout timescale is on the order of the orbital period, which is relatively fast (1.3 week at 0.1 au, 0.8 yr at 1 au, and 4.5 yr at 3 au around an 1.3 M☉ star), although the the presence of overabundant sub-µm grains suggests that such rapid removal is not taking place in this system, likely due to the presence of gas. HD 23514’s chromospheric activity index (Fang et al., 2018) indicates that its stellar wind should not be exceptionally strong, suggesting the drag force is dominated by the P-R drag with timescales where is the speed of light, is the gravitational constant, and is the stellar mass. Within 1 au, the P-R timescale is a few 100s yr in HD 23514.
@@
Collisional grinding operates in a cascade fashion among different sizes of particles with a loss time that depends on the properties of the particle swarm (belt location and width, total mass and the size of the particle) and stellar mass. Continuous grindings supply enough new grains against the loss from radiation blowout and P-R drag, therefore sustaining the dust level over the collisional timescales (). The collisional timescale can be estimated as formulated by Equ.(13) in Wyatt et al. 2007 where the largest unknown properties are the total disk mass (), the maximum size () of the planetesimals, and belt location ().
@@
Based on the SED modeling described above, we can set 3 au with a width = 0.1 for simplicity. If we take the SED derived dust mass (up to 1 mm) at the face value, is 500 yr for of 1 mm. However, if we extrapolate the dust mass to 1 km size planetesimals (i.e., 0.167 as the disk mass), 5 yr in this case. The collisional timescale is inversely proportional to the total disk mass if extrapolated to larger size of planetesimal (by a factor of ), and is directly related to the size of colliding bodies. If the disk in HD 23514 formed by a past giant impact at 3 au, it would remain infrared bright for 103-4 yr, consistent with the previous studies (Jackson & Wyatt, 2012).
@@
An alternative way to support the giant impact hypothesis for HD 23514 is to examine the mass loss rate and back track . From the SED model, we estimate the dust mass for the small grains, the ones that give rise to the solid-state feature, is g. These 1 µm grains are expected to be ejected from the system on the blowout timescale, which is 1 yr in HD 23514. Sustaining this high mass loss rate through collisional cascades would requires a total planetesimal mass of 1260 over the age of the star (150 Myr) (without counting the loss from P-R drag). This would require that HD 23514 has a 35 times more mass than the minimum mass solar nebula (MMSN) at birth (assuming 100:1 gas to dust mass ratio). It is then unlikely such a high rate has been sustained over the age of the system, i.e., the high mass loss rate is transient, consistent with the giant impact hypothesis.

### v2

@@ C.4 The Nature of the 15 µm bump in the Spitzer/IRS Spectrum
We investigated whether the 15 µm bump in the Spitzer spectrum as shown in Figure 3 could be due to data reduction artifacts by examining the CASSIS IRS spectra of a handful debris systems (e.g., HD 172555, HD 15407 and HD 69830, and others), which are brighter or at a similar brightness as HD 23514 and observed in the same observing mode, using the CASSIS products. We found no similar bump near 15 µm, suggesting that the 15 µm bump is likely astrophysical in nature.
@@
Appendix D Estimating the Dust Location, Blowout Sizes, and Various Timescales
D.1 Spectral Energy Distribution Models
@@
The purpose of the SED modeling is to estimate the disk extent and to illustrate that the bulk of the silica dust is colocated with the hot molecular gas (from the temperature argument). The derived parameters are then used to estimate various timescales that are relevant in HD 23514.
@@
Given the number of assumed parameters, the SED models are extremely degenerate and sensitive to the disk density distribution, and only serve as a zero-order estimate without other constraints like resolved disk images. We stress that the SED modeling results are only applicable to the part of the disk where its bulk emission is fully captured by the SED measurements. The assumptions for our SED models are (1) the star is the only heating source and in an optically thin environment, (2) there is one single, flat (constant surface density), continuous disk with a radial range of inner () and outer () radii from the star, and (3) the dust grains are all in one uniform size distribution with a size power-law index of 3.5 and a maximum size of 1000 µm. We explored several minimum grain sizes and found that a minimum size of 0.1–0.5 µm silica- and silicate-like grains are needed in order to produce the prominent 10/20 µm features, even though the typical radiation blowout size is 1 µm given the stellar mass and luminosity (1.3 and 3 , respectively). The dynamical stability of these small grains is subject to their exact optical properties in response to the stellar radiation, which is not explored/considered here. We consider that the impact of other parameters such as the maximum grain size () and surface density power-law is minimal in a way that (1) affects the total dust mass in a fashion under the assumed size distribution, and (2) a steeper surface density distribution would extend so enough colder material can account for the total emission.
@@
A large uncertainty in our SED model resides in the grain composition. The dust temperatures depend sensitively on the grain absorption efficiency at the wavelengths where the star emits most its energy (i.e., UV and optical for HD 23514). Submicron grains are generally hotter than their larger counterparts and the exact temperatures can be vastly different between different types of composition as shown in the left panel of Figure 9. Because the carbonaceous and silica grains are less absorptive than the typical silicates in the UV and optical, they generally have lower temperatures than those of silicates at the same stellocentric distance from the star.
@@
Determining detailed dust mineralogy to infer the exact dust composition is beyond the scope of this short paper. We only explored limited dust compositions that are commonly found in thermally processed dust in circumstellar environments. For the dust emission calculation, we adopted the grain properties computed by optool (Dominik et al., 2021) using the Distribution of Hollow Spheres (DHS, Min et al. 2005) approach with an irregularity factor of 0.8. We found that the system’s SED can be well represented by three types of dust: silica (density = 2.65 g cm-3, Henning & Mutschke 1997), pyroxene ( = 3.2 g cm-3, Dorschner et al. 1995) and amorphous carbon ( = 1.85 g cm-3, Zubko et al. 1996) where the first two give rise to the prominent dust features in the mid-IR while the amorphous carbon and a small fraction of Mg-rich pyroxene grains contribute the underlining featureless dust continuum as shown in Figure 4. A single component disk with = 0.1 au and 2–3 au gives a good fit to the measured SED except that the silica grains need to be restricted within 0.1–0.2 au to account for enough flux in the 3–5 µm range without changing the contrast/ratio between the 10 and 20 µm features (the right panel of Figure 9). In other words, having silica grains beyond 1 au would shift the near-IR flux toward the far-IR because of its low thermal temperature outside 1 au, resulting in poor fits between 3 and 5 µm. We note that the model produces a much narrower 10 µm feature than the observed one, likely due to the mismatch in the various silica polymorphs (Koike et al., 2013).
@@
Based on the SED model, the system’s IR fractional luminosity () is 1.76 with 70% coming from the featureless dust continuum. The total dust mass is g (1.67) and the feature-producing grains account for 5% of the mass. Because the SED models prefer a confined (sub-au) region for the silica grains, the total silica dust mass (integrated to 1 mm in size) is on the order of 1% (1022 g) of the total dust mass. In summary, the bulk of the dust in HD 23514 is within 3 au and the silica dust is mostly located inside the sub-au region, similar to the location of the hot molecular gas found in Appendix C.1. HD 23514 was not detected by the ALMA 1.3 mm shallow survey of the Pleiades (Sullivan et al., 2022), suggesting lack of cold (40 K) dust, although the ALMA data could not exclude a less dusty planetesimal belt like our current Kuiper belt. In conclusion, the HD 23514 system does not possess a massive (tens of au) Kuiper-belt-like analog that is found in 20% of solar-like stars (Sierchio et al., 2014; Matrà et al., 2025).
D.2 Radiation-pressure Blowout Sizes around HD 23514
@@
In gas-free debris disks, stellar radiation pressure is one of the most important forces acting on grains that counters with the star’s gravity and dictates the dust dynamics, which can be characterized by the parameter defined as the ratio of radiation pressure to gravitational forces:
@@
where and are, respectively, the luminosity and mass of the star, is the grain size, is the grain density, and is the radiation-pressure efficiency averaged over stellar spectrum (Burns et al., 1979). The last three parameters have strong dependency on the exact grain properties (composition, size, density, and porosity, as detailed in Arnold et al. 2019). Grains with get pushed into hyperbolic orbits and leave the system, while grains with are on elliptical orbits and spiral inward due to Poynting-Robertson (PR) drag. The radiation-pressure blowout size is defined as , and the exact size depends on the type of star, and most importantly, the grain properties as illustrated in Figure 10. The values are calculated using the stellar Kurucz model described in Appendix A.1, the grain compositions adopted in Appendix D.1 with two additions (olivine: 3.71 from Dorschner et al. 1995, and silica: 2.65 from Kitamura et al. 2007), and assuming compact (i.e., ideal with zero porosity) grain shape. For nominal silicate and carbonaceous compositions, grains smaller than the blowout size (1 µm for HD 23514) are ejected at a terminal velocity of (Su et al., 2005). While for silica composition, some (such as the one with the optical constant from Henning & Mutschke 1997) could have 0.5 for all sizes, and some (from Kitamura et al. 2007) would have 0.5 grains within two blowout sizes. This also illustrates that blowout size is not a strict number, highly depending on the exact composition. We stress the importance of grain optical properties in the UV and optical wavelengths where the star emits most of its energy in estimating the grain’s values.
@@ D.3 Relevant Timescales
For a dust-dominated debris disk, the dominant grain loss mechanisms are collisional grinding and PR/stellar wind drags. Once the grains generated by collisional cascades are smaller than the radiation blowout size (1 µm in HD 23514 when the ratio between the radiation pressure force and gravity, defined as , equal to 0.5), they are ejected from the system by radiation pressure under normal conditions. The blowout timescale is on the order of the orbital period, which is relatively fast (1.3 week at 0.1 au, 0.8 yr at 1 au, and 4.5 yr at 3 au around an 1.3 star), although the the presence of overabundant submicron grains suggests that such rapid removal is not taking place in this system, likely due to the presence of gas. HD 23514’s chromospheric activity index (Fang et al., 2018) indicates that its stellar wind should not be exceptionally strong, suggesting the drag force is dominated by the PR drag with timescale where is the speed of light, is the gravitational constant, and is the stellar mass. Within 1 au, the PR timescale is a few 100s yr in HD 23514.
@@
Collisional grinding operates in a cascade fashion among different sizes of particles with a loss time that depends on the properties of the particle swarm (belt location and width, total mass, and the size of the particle) and stellar mass. Continuous grindings supply enough new grains against the loss from radiation blowout and PR drag, therefore sustaining the dust level over the collisional timescale (). The collisional timescale can be estimated as formulated by Equation (13) in Wyatt et al. 2007 where the largest unknown properties are the total disk mass (), the maximum size () of the planetesimals, and belt location ().
@@
Based on the SED modeling described above, we can set 3 au with a width = 0.1 for simplicity. If we take the SED-derived dust mass (up to 1 mm) at face value, is 500 yr for of 1 mm. However, if we extrapolate the dust mass to 1 km size planetesimals (i.e., 0.167 as the disk mass), 5 yr in this case. The collisional timescale is inversely proportional to the total disk mass if extrapolated to larger sizes of planetesimals (by a factor of ), and is directly related to the sizes of colliding bodies. If the disk in HD 23514 formed by a past giant impact at 3 au, it would remain IR bright for 103-4 yr, consistent with previous studies (Jackson & Wyatt, 2012).
@@
An alternative way to support the giant-impact hypothesis for HD 23514 is to examine the mass-loss rate and backtrack . From the SED model, we estimate the dust mass for the small grains, the ones that give rise to the solid-state feature, is g. These 1 µm grains are expected to be ejected from the system on the blowout timescale, which is 1 yr in HD 23514. Sustaining this high mass-loss rate through collisional cascades would requires a total planetesimal mass of 1260 over the age of the star (150 Myr) without counting the loss from PR drag. This would require that the HD 23514 system is 35 times more massive than the MMSN at birth (assuming 100:1 gas-to-dust mass ratio). It is then unlikely such a high rate has been sustained over the age of the system, i.e., the high mass-loss rate is transient, consistent with the giant-impact hypothesis.